OpenStudy (anonymous):

Line AB has an equation of a line y =5x − 2. Which of the following could be an equation for a line that is parallel to line AB?

OpenStudy (anonymous):

y = 5x + 3 y = 1 over 5x + 3 y = −5x + 3 y = −1 over 5x + 3

OpenStudy (anonymous):

is it a because its following the y=mx+b format?

OpenStudy (mrnood):

in the y=mx+b format the 'm' tells oyu the slope parallel lines have the SAME slope (the 'b' value does not affect whether they are parallel) So look at all the answers and find one that has the same value for m as y =5x − 2

OpenStudy (anonymous):

is it c?

OpenStudy (mrnood):

the original equation is y =5x − 2 c is y = -5x +3 in the format y = mx+b do they have the SAME value for m?

OpenStudy (anonymous):

im so confused

OpenStudy (mrnood):

with what?

OpenStudy (anonymous):

with like the format

OpenStudy (anonymous):

cause you said find the one that has the same value and none of them do

OpenStudy (mrnood):

look at the original equation y =5x − 2 now look at the format y = mx + b what is the value of m in the original equation?? what is the value for b in the original equation?

OpenStudy (mrnood):